Solve the following polynomial equation using formula method

3x²+9x+7=0


Solve the following polynomial equation using factorization method

X³+2x²‐9x‐18=0


1
Expert's answer
2021-12-06T17:13:57-0500

Let us solve the polynomial equation "3x^2+9x+7=0" using formula method. It follows that "D=9^2-4\cdot 3\cdot 7=81-84=-3." Therefore, the equation has no real solutions. The complex solutions are the following:

"x_1=\frac{-9+i\sqrt{3}}{6}" and "x_2=\frac{-9-i\sqrt{3}}{6}."


Let us solve the polynomial equation "x^3+2x^2-9x-18=0" using factorization method. It follows that

"x^3+2x^2-9x-18=x^2(x+2)-9(x+2)=(x^2-9)(x+2)=(x-3)(x+3)(x+2),"

and hence the equation "x^3+2x^2-9x-18=0" is equivalent to "(x-3)(x+3)(x+2)=0."

Therefore, it has the following solutions:

"x_1=3,x_2=-3" and "x_3=-2."